OpenStudy (anonymous):

A tennis ball is thrown straight up with an inital speed of 22.5 m/s. It is caught at the same distance abovethe ground. How high doe the ball rise? Please show work.

OpenStudy (anonymous):

Max height is when the vertical speed reaches zero. You can calculate the time it took to get to the top by working out t 9.8t = 22.5 With that you can solve d = 22.5t - 4.9t^2

OpenStudy (doc.brown):

\[v^2=v_0^2+2a(y-y_0)\]\[0=v_0^2+2a(y-0)\]\[-v_0^2=2ay\]\[y=\frac{-v_0^2}{2a}\]